a snake's length ranges from 48 to 130 inches. find an absolute value inequality equation to show the range of the snake's length

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

130[=<(x)=>]48 where x=the snake's length.

Answer by stanbon(48546)

a snake's length ranges from 48 to 130 inches. find an absolute value inequality equation to show the range of the snake's length
-----------------
Find the mid-point: (48+130)/2 = 24+65 = 89
------------------
Find the distance from 89 to an end-point: 89-48 = 41
-------------------
Your Problem:
|x-89|<=41
================
Cheers,
Stan H.